OpenStudy (anonymous):

Can someone help me or show me how to do this problem? " What is the mass in grams of each of the following samples? 1.505 mol of Ti? " Thank you!!

OpenStudy (anonymous):

first molar mass of Ti is about 47.8 gr/mol. then we know that we have 1.505 mol of Ti. so we use this formula \[n=\frac{ m }{M }\]n=1.505 M=47.8 gr/mol m=? gr \[m=1.505 \times 47.8 \approx 72\]

OpenStudy (anonymous):

we know that \[n = \frac{ m }{ M }\] with : ' n ' in moles and ' m ' in grams and ' M ' is nbr atomique so we have \[m = n * M = 1.505 * 47.8 = 72 g\]